# Is Long Context All You Need? Leveraging LLM’s Extended Context for NL2SQL Experiments, Analysis and Benchmark Paper

### v5

@@ Abstract.
PVLDB, 14(1): XXX-XXX, 2020.
doi:XX.XX/XXX.XX This work is licensed under the Creative Commons BY-NC-ND 4.0 International License. Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license. For any use beyond those covered by this license, obtain permission by emailing info@vldb.org. Copyright is held by the owner/author(s). Publication rights licensed to the VLDB Endowment.
Proceedings of the VLDB Endowment, Vol. 14, No. 1 ISSN 2150-8097.
doi:XX.XX/XXX.XX
@@ 1. Introduction
Recent advancements in LLMs have particularly focused on enhancing their retrieval and reasoning abilities and expanding their context windows, thus broadening the scope of their potential applications. The ability to process and retain longer sequences of information within the context window empowers LLMs to capture nuanced dependencies and intricate relationships within the input data, offering unprecedented possibilities for improved language understanding and generation. One domain that stands to significantly benefit from these advancements is Natural Language to SQL (NL2SQL). NL2SQL is a challenging task that entails translating natural language questions into structured SQL queries that can be executed on a database. The inherent ambiguity of natural language questions, coupled with the necessity for a deep understanding of complex database schemas and semantics, makes NL2SQL a very challenging problem (Floratou et al., 2024). Recent works (Pourreza et al., 2024; Maamari et al., 2024; Caferoğlu and Ulusoy, 2024; Gao et al., 2024) created NL2SQL pipelines involving schema linking, self-consistency, and self-correction with multiple LLM calls. These solutions carefully create prompts that include various context information, and use Chain-of-Thought (CoT)(Liu and Tan, 2023; Zhou et al., 2022) reasoning and/or in-context-learning.
@@
CHASE-SQL1112nd on the leaderboard as of December 17th, 2024. Long context NL2SQL pipeline was used as one of the NL2SQL pipelines to generate multiple output candidates (Pourreza et al., 2024) 74.46
@@
Self-consistency is another popular technique used in state-of-the-art NL2SQL systems (Wang et al., 2022; Talaei et al., 2024; Pourreza et al., 2024; Dong et al., 2023; Gao et al., 2023; Gao et al., 2024). The idea is to generate multiple output candidates and pick the most likely one according to some rules (e.g., majority voting) or a fine-tuned picker model. This works nicely due to the stochastic nature of LLMs, where different models (and even the same model) generate varying outputs for the same input. Self-consistency is orthogonal to our work, and one can combine the two (Pourreza et al., 2024) for higher accuracy (see Section 6.2 for more discussion). In this work we focus on identifying and providing additional contextual information for ICL, leveraging the extended context window of a long-context LLM.
@@ 2. Related Works
One can refine the output through multiple LLM calls, verifying, fixing and rewriting the SQL outputs. It is also common to sample multiple answer candidates and select (self-consistency) the most probable or consistent output to improve the quality (Wang et al., 2022; Talaei et al., 2024; Pourreza et al., 2024; Dong et al., 2023; Gao et al., 2023; Gao et al., 2024). In this work, we focus on leveraging the long context ICL for a typical NL2SQL pipeline – which is orthogonal and can be used with self-consistency. (Pourreza et al., 2024) uses our long context pipeline to generate candidates along with a fine-tuned picker.
@@ 3.1. Generate
Synthetic examples for many-shot ICL: While prior works on ICL for NL2SQL focused on selecting a few handful of examples (3-5) for ICL (Nan et al., 2023; Pourreza and Rafiei, 2024), we generate many examples for many-shot ICL. Many-shot ICL is shown to improve LLM generation quality on many different domains (Agarwal et al., 2024). Here we generate and use tens and hundreds of examples with the extended context window. The examples are also generated at inference time, as prescribed in (Pourreza et al., 2024). The online generation allows bypassing the example retrieval process and generate examples using schema elements more relevant to the user question (we pass the entire schema for a given database along with LLM-based relevant column selection results). The model generates synthetic examples including both the output SQL and input natural language question. The generated SQL queries follow the target benchmark datasets with respect to the common SQL features The structural similarities can help guide the model to generate with similar SQL structures and features.
@@ 3.3. Verify
We use the untuned gemini-1.5-pro LLM to verify the correctness of the final output. The model is given the entire database table schema and question (with additional hints, if provided) to judge. We believe that fine-tuning a verifier or a picker for self-consistency(Wang et al., 2022; Pourreza et al., 2024; Zheng et al., 2024) can further improve accuracy. While we focus on leveraging the long context ICL for a typical NL2SQL pipeline with simpler techniques (self-correction, verification without fine-tuning), the aforementioned techniques are orthogonal and can be used in conjunction. We demonstrate how this verification step can impact the final performance and the generation cost in Section 5.
@@ 5. Ablation Study
Because the latest long-context LLM, gemini-pro-1.5, has strong retrieval capabilities, we observe overall performance improvements as more contextual information is included. However, when we categorize BIRD dev questions by difficulty level, we find that different types of contextual information benefit different types of queries. For instance, providing extra sample column values (or providing them all for disambiguation) is more helpful for challenging questions, where the question is more nuanced and/or ambiguous - allowing the model to leverage exact column values to choose the correct column and join paths. Synthetic examples show the opposite where it helps with simple and moderate questions, but hurt the challenging questions. Synthetic examples are generated to illustrate common SQL features and clauses using the schema elements from the target DB (Pourreza et al., 2024). They tend to be simpler and involve less ambiguous mappings between natural language questions and SQL queries. The hints are also interesting in that it appears to be one of the most critical ingredients for accurate SQL generation and it benefits moderate questions the most. BIRD datasets contain hints as to clarify nuanced concepts (e.g., “eligible free rate for K-12”) with a correct expression (“Eligible free rate for K-12 = ‘Free Meal Count (K-12)‘ / ‘Enrollment (K-12)‘”), and moderate queries often consist of more nuanced concepts that require such clarifications.
@@
Table 9 reports the mini-ablation study, which analyzes the efficacy of the most controversial (expensive) long context techniques - disambiguation and many-shot ICL with synthetic examples - across other benchmark datasets. Running the full pipeline without each of the aforementioned techniques reduces both accuracy and latency (or context size). The performance implications of skipping disambiguation (w/o disambiguation) are limited on SPIDER and KaggleDBQA benchmarks, but much more pronounced on the BEAVER dataset. This is because disambiguation is only triggered for executable SQL queries that return empty result sets. Such invalid predictions happens more frequently in the BEAVER data warehouse (dw) dataset, as its schema is significantly more complex and requires more joins (4.25 on the average) compared to other datasets. Similarly, skipping many-shot ICL with synthetic examples (w/o synthetic examples) reduces both accuracy and latency, but the impact is limited on SPIDER and KaggleDBQA. On the contrary, w/o synthetic examples on BEAVER does not change the accuracy, but also increasing the latency significantly due to more frequent disambiguation attempts. This suggests that, overall, the generated examples are more misleading than helpful, but resulted in fewer empty results. This is expected, as the synthetic example generation primarily targets simpler query structures (Pourreza et al., 2024) rather than star-schema (data warehouse) queries involving many joins. It is also interesting to note that disambiguation and synthetic examples have very different performance characteristics. As shown in the full ablation study with BIRD dev (Table 8), disambiguation is more helpful for challenging questions, whereas synthetic examples contribute more to solving simpler problems. We observe the same trends across the other benchmark datasets, as illustrated in Table 10.
@@ 6.2. Further performance improvement via self-consistency
Self-consistency is another very popular technique for NL2SQL (Wang et al., 2022; Talaei et al., 2024; Pourreza et al., 2024; Dong et al., 2023; Gao et al., 2023; Gao et al., 2024), where multiple candidates are generated, and the most consistent answer or the best one is selected by a fine-tuned picker. As demonstrated by the recent state-of-the-art approaches (Pourreza et al., 2024; Gao et al., 2024) on the BIRD leaderboard, self-consistency has become a crucial technique for achieving high accuracy in NL2SQL. This involves generating multiple candidates using different generators (a.k.a. NL2SQL pipelines) and each generator trying multiple times, and choosing the best one. One of the state-of-the-arts, CHASE-SQL, in Table 1 uses our long context pipeline as one of the three candidate generators, contributing to its top-tier performance on the BIRD leaderboard. One caveat of self-consistency is that it can quickly become expensive in terms of latency and the number of LLM calls. Self-consistency is orthogonal to our work, and can further improve the NL2SQL performance if used together with our pipeline. Table 11 illustrates the impact of generating multiple output candidates from our long context pipeline using a generation temperature of 0.5. With an oracle picker (i.e., at least one of the candidates is correct per question) we can uplift the accuracy beyond 70%. In (Pourreza et al., 2024; Gao et al., 2024), the candidate sets are generated using multiple pipelines, as different strategies result in more diverse outputs, further boosting the final accuracy of self-consistency.

### v6

@@ Abstract.
PVLDB, 18(8): 2735 - 2747, 2025.
doi:10.14778/3742728.3742761 This work is licensed under the Creative Commons BY-NC-ND 4.0 International License. Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license. For any use beyond those covered by this license, obtain permission by emailing info@vldb.org. Copyright is held by the owner/author(s). Publication rights licensed to the VLDB Endowment.
Proceedings of the VLDB Endowment, Vol. 18, No. 8 ISSN 2150-8097.
doi:10.14778/3742728.3742761
@@ 1. Introduction
Recent advancements in LLMs have particularly focused on enhancing their retrieval and reasoning abilities and expanding their context windows, thus broadening the scope of their potential applications. The ability to process and retain longer sequences of information within the context window empowers LLMs to capture nuanced dependencies and intricate relationships within the input data, offering unprecedented possibilities for improved language understanding and generation. One domain that stands to significantly benefit from these advancements is Natural Language to SQL (NL2SQL). NL2SQL is a challenging task that entails translating natural language questions into structured SQL queries that can be executed on a database. The inherent ambiguity of natural language questions, coupled with the necessity for a deep understanding of complex database schemas and semantics, makes NL2SQL a very challenging problem (Floratou et al., 2024). Recent works (Pourreza et al., 2025; Maamari et al., 2024; Caferoğlu and Ulusoy, 2024; Gao et al., 2024) created NL2SQL pipelines involving schema linking, self-consistency, and self-correction with multiple LLM calls. These solutions carefully create prompts that include various context information, and use Chain-of-Thought (CoT)(Liu and Tan, 2023; Zhou et al., 2022) reasoning and/or in-context-learning.
@@
CHASE-SQL1112nd on the leaderboard as of December 17th, 2024. Long context NL2SQL pipeline was used as one of the NL2SQL pipelines to generate multiple output candidates (Pourreza et al., 2025) 74.46
@@
Self-consistency is another popular technique used in state-of-the-art NL2SQL systems (Wang et al., 2022; Talaei et al., 2024; Pourreza et al., 2025; Dong et al., 2023; Gao et al., 2023; Gao et al., 2024). The idea is to generate multiple output candidates and pick the most likely one according to some rules (e.g., majority voting) or a fine-tuned picker model. This works nicely due to the stochastic nature of LLMs, where different models (and even the same model) generate varying outputs for the same input. Self-consistency is orthogonal to our work, and one can combine the two (Pourreza et al., 2025) for higher accuracy (see Section 6.2 for more discussion). In this work we focus on identifying and providing additional contextual information for ICL, leveraging the extended context window of a long-context LLM.
@@ 2. Related Works
One can refine the output through multiple LLM calls, verifying, fixing and rewriting the SQL outputs. It is also common to sample multiple answer candidates and select (self-consistency) the most probable or consistent output to improve the quality (Wang et al., 2022; Talaei et al., 2024; Pourreza et al., 2025; Dong et al., 2023; Gao et al., 2023; Gao et al., 2024). In this work, we focus on leveraging the long context ICL for a typical NL2SQL pipeline – which is orthogonal and can be used with self-consistency. (Pourreza et al., 2025) uses our long context pipeline to generate candidates along with a fine-tuned picker.
@@ 3.1. Generate
Synthetic examples for many-shot ICL: While prior works on ICL for NL2SQL focused on selecting a few handful of examples (3-5) for ICL (Nan et al., 2023; Pourreza and Rafiei, 2024), we generate many examples for many-shot ICL. Many-shot ICL is shown to improve LLM generation quality on many different domains (Agarwal et al., 2024). Here we generate and use tens and hundreds of examples with the extended context window. The examples are also generated at inference time, as prescribed in (Pourreza et al., 2025). The online generation allows bypassing the example retrieval process and generate examples using schema elements more relevant to the user question (we pass the entire schema for a given database along with LLM-based relevant column selection results). The model generates synthetic examples including both the output SQL and input natural language question. The generated SQL queries follow the target benchmark datasets with respect to the common SQL features The structural similarities can help guide the model to generate with similar SQL structures and features.
@@ 3.3. Verify
We use the untuned gemini-1.5-pro LLM to verify the correctness of the final output. The model is given the entire database table schema and question (with additional hints, if provided) to judge. We believe that fine-tuning a verifier or a picker for self-consistency(Wang et al., 2022; Pourreza et al., 2025; Zheng et al., 2024) can further improve accuracy. While we focus on leveraging the long context ICL for a typical NL2SQL pipeline with simpler techniques (self-correction, verification without fine-tuning), the aforementioned techniques are orthogonal and can be used in conjunction. We demonstrate how this verification step can impact the final performance and the generation cost in Section 5.
@@ 5. Ablation Study
Because the latest long-context LLM, gemini-pro-1.5, has strong retrieval capabilities, we observe overall performance improvements as more contextual information is included. However, when we categorize BIRD dev questions by difficulty level, we find that different types of contextual information benefit different types of queries. For instance, providing extra sample column values (or providing them all for disambiguation) is more helpful for challenging questions, where the question is more nuanced and/or ambiguous - allowing the model to leverage exact column values to choose the correct column and join paths. Synthetic examples show the opposite where it helps with simple and moderate questions, but hurt the challenging questions. Synthetic examples are generated to illustrate common SQL features and clauses using the schema elements from the target DB (Pourreza et al., 2025). They tend to be simpler and involve less ambiguous mappings between natural language questions and SQL queries. The hints are also interesting in that it appears to be one of the most critical ingredients for accurate SQL generation and it benefits moderate questions the most. BIRD datasets contain hints as to clarify nuanced concepts (e.g., “eligible free rate for K-12”) with a correct expression (“Eligible free rate for K-12 = ‘Free Meal Count (K-12)‘ / ‘Enrollment (K-12)‘”), and moderate queries often consist of more nuanced concepts that require such clarifications.
@@
Table 9 reports the mini-ablation study, which analyzes the efficacy of the most controversial (expensive) long context techniques - disambiguation and many-shot ICL with synthetic examples - across other benchmark datasets. Running the full pipeline without each of the aforementioned techniques reduces both accuracy and latency (or context size). The performance implications of skipping disambiguation (w/o disambiguation) are limited on SPIDER and KaggleDBQA benchmarks, but much more pronounced on the BEAVER dataset. This is because disambiguation is only triggered for executable SQL queries that return empty result sets. Such invalid predictions happens more frequently in the BEAVER data warehouse (dw) dataset, as its schema is significantly more complex and requires more joins (4.25 on the average) compared to other datasets. Similarly, skipping many-shot ICL with synthetic examples (w/o synthetic examples) reduces both accuracy and latency, but the impact is limited on SPIDER and KaggleDBQA. On the contrary, w/o synthetic examples on BEAVER does not change the accuracy, but also increasing the latency significantly due to more frequent disambiguation attempts. This suggests that, overall, the generated examples are more misleading than helpful, but resulted in fewer empty results. This is expected, as the synthetic example generation primarily targets simpler query structures (Pourreza et al., 2025) rather than star-schema (data warehouse) queries involving many joins. It is also interesting to note that disambiguation and synthetic examples have very different performance characteristics. As shown in the full ablation study with BIRD dev (Table 8), disambiguation is more helpful for challenging questions, whereas synthetic examples contribute more to solving simpler problems. We observe the same trends across the other benchmark datasets, as illustrated in Table 10.
@@ 6.2. Further performance improvement via self-consistency
Self-consistency is another very popular technique for NL2SQL (Wang et al., 2022; Talaei et al., 2024; Pourreza et al., 2025; Dong et al., 2023; Gao et al., 2023; Gao et al., 2024), where multiple candidates are generated, and the most consistent answer or the best one is selected by a fine-tuned picker. As demonstrated by the recent state-of-the-art approaches (Pourreza et al., 2025; Gao et al., 2024) on the BIRD leaderboard, self-consistency has become a crucial technique for achieving high accuracy in NL2SQL. This involves generating multiple candidates using different generators (a.k.a. NL2SQL pipelines) and each generator trying multiple times, and choosing the best one. One of the state-of-the-arts, CHASE-SQL, in Table 1 uses our long context pipeline as one of the three candidate generators, contributing to its top-tier performance on the BIRD leaderboard. One caveat of self-consistency is that it can quickly become expensive in terms of latency and the number of LLM calls. Self-consistency is orthogonal to our work, and can further improve the NL2SQL performance if used together with our pipeline. Table 11 illustrates the impact of generating multiple output candidates from our long context pipeline using a generation temperature of 0.5. With an oracle picker (i.e., at least one of the candidates is correct per question) we can uplift the accuracy beyond 70%. In (Pourreza et al., 2025; Gao et al., 2024), the candidate sets are generated using multiple pipelines, as different strategies result in more diverse outputs, further boosting the final accuracy of self-consistency.